FMA Inc has issued a $1000 par value bond that matures in 14 years. The bond pays semi-annual coupons at a rate of 7.5% APR compounded semi-annually, with first coupon payment due 6-months from today. What is the bond's price if the market requires a 9.5% yield to maturity on this bond?
Answer #1

Par Value =1000
Number of Years =14*2 =28
Semi annual Coupon =7.5%*1000/2 =37.50
Semi annual YTM =9.5%/2 =4.75%
Bond's Price =PV of Coupons +PV of Par Value =37.50*((1-(1+4.75%)^-28)/4.75%) +1000/(1+4.75%)^28=846.88